OpenStudy (anonymous):

what is the inverse of 2(sqaure root)-3x - 8

OpenStudy (anonymous):

2\[2\sqrt{3x}-8\]

OpenStudy (jdoe0001):

to find the inverse "relation" of the function... firstly swap about the variables \(\bf {\color{brown}{ y}}=2\sqrt{3{\color{blue}{ x}}}-8\qquad inverse\implies {\color{blue}{ x}}=2\sqrt{3{\color{brown}{ y}}}-8\impliedby f^{-1}\) then solve for "y"